Find the dimensions of the rectangle whose width is 8cm less than its length and whose area is 20 cm^2.
Answer by CubeyThePenguin(3113)

area = width * length

20 = (L)(L - 8)
20 = L^2 - 8L
0 = L^2 - 8L - 20
0 = (L - 10)(L + 2)

Distance is positive, so the rectangle is 10 cm long and 2 cm wide.